# Model Context Protocol (MCP) Tool Descriptions Are Smelly! Towards Improving AI Agent Efficiency with Augmented MCP Tool Descriptions

### v1

@@ 6. Implications
In this section, we discuss the implications of our findings for major stakeholder groups: (i) MCP developers, (ii) ecosystem maintainers (e.g., protocol and registry maintainers), (iii) MCP users (e.g., agent developers), and (iv) researchers.
@@ 6.2. Implications for Ecosystem Maintainers
Registry maintainers should integrate rubric-based smell detection and quality scoring as built-in services across MCP registries and marketplaces. As RQ-1 shows, smells are pervasive across both official and community servers, indicating that no maintainer group consistently produces high-quality descriptions. This finding implies that quality cannot be reliably delegated to individual teams. MCP registries, such as Smithery, Glama, or Cloudflare Workers, serve as centralized hubs for discovering and deploying MCP servers; however, their existing review processes primarily focus on infrastructure setup or package-level checks, and rarely assess the quality of tool descriptions. Ecosystem maintainers can respond to the pervasive smells by running FM-based scans over published servers, with smell summaries and augmentation-aware quality badges, and issuing warnings when a tool or server falls below agreed thresholds. Because the FM-based scanner shows stable agreement across models, these diagnostics can function similarly to security advisories in registries such as npm, encouraging developers to submit higher-quality descriptions or maintain multiple augmented profiles to reduce integration friction.
@@ 7.1. External validity
While we detect and optimize smells across 856 tools from 103 MCP servers, performance evaluation is conducted only on the subset of tools and servers included in the MCP-Universe benchmark. Specifically, MCP-Universe covers 202 tools drawn from 18 MCP servers, which represent a strict subset of the full corpus analyzed in RQ1 and RQ2. In addition, due to the high computational cost of running the full benchmark, we do not evaluate all models included in MCP-Universe. To mitigate these limitations, we adopt MCP-Universe because it inherently covers a diverse set of tasks spanning multiple domains, MCP servers, and evaluators, enabling robust execution-based assessment without additional sampling. We also intentionally select a heterogeneous model set, covering a frontier proprietary model (GPT-4.1), large open-weight models (Qwen3-Coder-480B-A35B and GLM-4.5), and a smaller-sized model (Qwen3-Next-80B-A3B-Instruct), to capture variation across model families while balancing evaluation cost.
The evaluation of the fully augmented tool description is constrained by context window limits of the models and by the verbosity of some tools. For Qwen3-Next-80B-A3B-Instruct, we exclude the Parameter Explanation and Examples components to avoid context overflow. Similarly, for the Browser Automation domain, the excessive length of execution examples necessitates excluding examples for all models. As a result, the evaluation for Qwen3-Next-80B-A3B-Instruct reflects a partial augmentation setting, and the full benefits of the proposed pipeline may not be realizable on resource-constrained models or highly verbose tool domains. We justify our choices by prioritizing components that are not otherwise available to the FM. Specifically, while Parameter Explanation is omitted, the MCP protocol still provides the input schema to the FM, partially compensating for its absence. In addition, our ablation study in RQ-3 shows that excluding Examples has a limited impact on performance in most settings, supporting this design choice under context constraints.
@@ 7.3. Internal validity
The original MCP-Universe study uses a SERP API-based Google Search MCP server with strict query limits. To ensure the uninterrupted execution of Web Searching tasks, we utilize an alternative Google Search MCP server. Differences in the underlying API platform may impact absolute performance in the Web Searching domain, regardless of the tool’s description quality. They may also explain the lack of performance improvement observed across all three MCP-Universe models. To mitigate this issue, we evaluate the Web Searching domain for the newly introduced model, Qwen3-Next-80B-A3B-Instruct, using the same Google Search MCP server for both the baseline and augmented tool descriptions. As a result, the statistical analyses for this model control for instrumentation differences, ensuring that observed performance changes are attributable to description augmenting rather than search API variability.
@@ 8. Conclusion
This paper presents the first systematic, large-scale evaluation of the quality of tool descriptions in the Model Context Protocol (MCP) ecosystem and their impact on FM-based agents. By analyzing 856 tools from both official and community-maintained servers and by conducting rubric-guided augmentation, benchmarking, and controlled studies, we establish tool descriptions as a critical but under-engineered artifact of agentic systems.
@@
ChatGPT-5.1, Anthropic Opus-4.5, and Gemini-3 Pro-Preview were used only for copy-editing and table formatting, in compliance with IEEE and ACM policies on AI use in publications.

### v2

@@ 6. Implications
This section outlines the practical and research implications of our results for four key audiences: (i) MCP developers, (ii) ecosystem maintainers (e.g., protocol and registry maintainers), (iii) MCP users (e.g., agent developers), and (iv) researchers.
@@ 6.2. Implications for Ecosystem Maintainers
Registry maintainers should integrate rubric-based smell detection and quality scoring as built-in services across MCP registries and marketplaces. As RQ-1 shows, smells are pervasive across both official and community servers, indicating that no maintainer group consistently produces high-quality descriptions. This finding implies that quality cannot be reliably delegated to individual teams. MCP registries, such as Glama, Smithery, or Cloudflare Workers, existing review processes primarily focus on infrastructure setup or package-level checks, and rarely assess the quality of tool descriptions. Ecosystem maintainers can respond to the pervasive smells by running FM-based scans over published servers, with smell summaries and augmentation-aware quality badges, and issuing warnings when a tool or server falls below agreed thresholds. Because the FM-based scanner shows stable agreement across models, these diagnostics can function similarly to security advisories in registries such as npm, encouraging developers to submit higher-quality descriptions or maintain multiple augmented profiles to reduce integration friction.
@@ 7.1. External validity
While we detect and optimize smells across 856 tools from 103 MCP servers, performance evaluation is conducted only on the subset of tools and servers included in the MCP-Universe benchmark. Specifically, MCP-Universe covers 202 tools drawn from 18 MCP servers, which represent a strict subset of the full corpus analyzed in RQ1 and RQ2. In addition, due to the high token cost of running the full benchmark, we do not evaluate all models included in MCP-Universe. To mitigate these limitations, we adopt MCP-Universe because it inherently covers a diverse set of tasks spanning multiple domains, MCP servers, and evaluators, enabling robust execution-based assessment without additional sampling. We also intentionally select a heterogeneous model set, covering a frontier proprietary model (GPT-4.1), large open-weight models (Qwen3-Coder-480B-A35B and GLM-4.5), and a smaller-sized model (Qwen3-Next-80B-A3B-Instruct), to capture variation across model families while balancing evaluation cost.
The feasibility of evaluating the fully augmented tool description depends on the available context window of the models and the verbosity of the tool descriptions. For Qwen3-Next-80B-A3B-Instruct, we exclude the Parameter Explanation and Examples components to avoid context overflow. Similarly, for the Browser Automation domain, the excessive length of execution examples necessitates excluding examples for all models. As a result, the evaluation for Qwen3-Next-80B-A3B-Instruct reflects a partial augmentation setting, and the full benefits of the proposed pipeline may not be realizable on resource-constrained models or highly verbose tool domains. We justify our choices by prioritizing components that are not otherwise available to the FM. Specifically, while Parameter Explanation is omitted, the MCP protocol still provides the input schema to the FM, partially compensating for its absence. In addition, our ablation study in RQ-3 shows that excluding Examples has a limited impact on performance in most settings, supporting this design choice under context constraints.
@@ 7.3. Internal validity
The original MCP-Universe study uses a SERP API-based Google Search MCP server with strict query limits. To ensure the uninterrupted execution of Web Searching tasks, we utilize an alternative Google Search MCP server. Differences in the underlying API platform may impact absolute performance in the Web Searching domain, regardless of the tool’s description quality. They may also explain the lack of performance improvement observed across all three MCP-Universe models. To mitigate this issue, we evaluate the Web Searching domain for the newly introduced model, Qwen3-Next-80B-A3B-Instruct, using the same Google Search MCP server for both the baseline and augmented tool descriptions. As a result, the statistical analyses for this model control for instrumentation differences, ensuring that measured differences in performance reflect description augmenting effects, as opposed to search API variability.
@@ 8. Conclusion
This research provides an extensive empirical characterization of MCP tool descriptions and evaluates the extent to which these descriptions affect FM-based agent execution. By analyzing 856 tools from both official and community-maintained servers and by conducting rubric-guided augmentation, benchmarking, and controlled studies, we establish tool descriptions as a critical but under-engineered artifact of agentic systems.
@@
We employed ChatGPT-5.2, Anthropic Opus-4.5, and Gemini-3 Pro-Preview solely for copy-editing and table layout preparation, consistent with IEEE and ACM policies governing AI use in scholarly work.